x mmol of KIO 3 reacts completely with y mmol of KI to give I 2 quantitatively. If z mmol of hypo are required for complete titration against this I 2 then which relation is not correct ? ​ a. z = 6x ​ b. 6y = 5z ​ c. x = 5y ​ d. x + y = 2z Please explain in detail.............

x mmol of KIO3 reacts completely with y mmol of KI to give I2 quantitatively. If z mmol of hypo are required for complete titration against this I2 then which relation is not correct ?
​ a.  z = 6x
​ b.  6y = 5z
​ c.  x = 5y
​ d.  x + y = 2z
Please explain in detail.............

Grade:12th pass

4 Answers

Vikas TU
4 years ago
Hiii Rather
From my point of view the option D should be not the ans.
Option D should be x+y = z not x+y = 2z
If this option will correct then both c and d would be correct.
RATHER YASIR
4 years ago
Hii Vikas sir......................................
Please Sir...Can you explain in detail.................................................................................
 
MANTHAN RAO
3 years ago
Correct option : (c)
Explanation:
  Here, 
KIO3 + 5KI + 3H2SO4 → 3I2 + 3H2O + 3K2SO4 
And 
6Na2S2O3 + 3I2 → 3Na2S4O6 + 6NaI
From the above two chemical equation, 
x = 1
, y = 5
and  z =6
Now, 
KIO3 + 6Na2S2O3 + 3H2SO4 → 3NaS4O6 + KI + 3H2O + 3Na2SO4 
We have , Coa (Concentration of oxidising agent) x Voa (volume of oxidising agent)/ 1 = Cra x Vra (Reducing agent) / 6 = > x/1 = z/6 [v = 1 (assumed)] => z = 6x
Similarly, when we will have KI as the oxidising agent and Na2S2O3 as the reducing agent : y/5 = z/6
 => 6y = 5z Now, x + y = (1 + 5) = 6
 = z so, option (c) 
 
 
MANTHAN RAO
3 years ago
 
Correct option : (c)
Explanation:
  Here, 
KIO3 + 5KI + 3H2SO4 → 3I2 + 3H2O + 3K2SO4 
And 
6Na2S2O3 + 3I2 → 3Na2S4O6 + 6NaI
From the above two chemical equation, 
x = 1
, y = 5
and  z =6
Now, 
KIO3 + 6Na2S2O3 + 3H2SO4 → 3NaS4O6 + KI + 3H2O + 3Na2SO4 
We have , Coa (Concentration of oxidising agent) x Voa (volume of oxidising agent)/ 1 = Cra x Vra (Reducing agent) / 6 = > x/1 = z/6 [v = 1 (assumed)] => z = 6x
Similarly, when we will have KI as the oxidising agent and Na2S2O3 as the reducing agent : y/5 = z/6
 => 6y = 5z Now, x + y = (1 + 5) = 6
 = z so, option (c) Option D should be x+y = z not x+y = 2z
IF YOUR OPTION IS CORRECT THAN THERE WILL BE TWO CORRECT ANSWER FOR THIS QUESTION THAT ARE (C) AND (D)